# Lessons from pendulums: A design comparison of three lab activities

### v1

@@ Abstract
We present three versions of a pendulum lab activity and explore how shared theoretical commitments, motivations, and aspirations can lead to divergences in curriculum design. Building on Boudreaux & Elby [boudreaux_how_2020] we provide another demonstration of how theoretical commitments and perspectives influence curriculum development. In this paper, the variations across these lab activities despite our shared theoretical bases constitutes the phenomena of interest. We argue that three main factors lead to the variations in our design decisions: different expectations and understandings of our student populations, variations in our ancillary pedagogical goals, and fine grained differences in our theoretical perspectives. By focusing on the differences between our labs, we highlight the limitations of theory alone to determine curriculum development and demonstrate the process of theory, in conversation with local particularities, guiding design.
@@ I Introduction
For longer than any of us has been learning physics, let alone teaching, curricula have included pendulums in introductory labs. Often, the goals have included students’ finding that the period of a pendulum has a small but measurable dependence on the amplitude of its swing, a result at odds with the simple harmonic oscillator model and Galileo’s claim in Two New Sciences [TwoNewSciences]. The lab provides a context for learning practices and values of empirical science: To see the dependence, students need both to take care in their measurements and to be aware of the precision they have obtained [^1] They do not need sophisticated technology! Mechanical stopwatches work fine..
We have all assigned pendulum labs at the start of our introductory courses, at our respective institutions, with goals of introducing practices and values of empirical science. The activities should promote students’ attention to experimental error as well as reflection and conversation about how science produces and assesses knowledge. These goals reflect our shared views of students: that they frame what is taking place based on what they have experienced, and what they have experienced in science labs has consisted mostly of step-by-step instructions for experiments that demonstrate or validate theoretical claims from lecture.
The similarities in our interests, goals, and views motivated some of us (Holmes, Hammer, Scherr, and Descamps) to collaborate in research on how students may shift in their framing during our labs, to be less about “doing school” and more about “doing physics” [smith_surprise_2018, smith_how_2020, smith_lets_2020]. That is, we have asked: when and how do students come to frame introductory labs in ways that support progress toward practices and values of empirical science? We were supported by the National Science Foundation for three years to collect and study data from our courses. [^2] NSF DUE Awards: 2000394, 2000739, 2000711. We often paid particular attention to the pendulum labs, because they were the students’ first engagement with activities different from their expectations. The project led to several publications analyzing student framings, including with respect to confirmation bias and epistemic agency [sundstrom_problematizing_2020, phillips_not_2021, jeon_2023, sundstrom_instructing_2023, descamps_dynamics_2024].
Throughout, we were aware of differences in our respective designs. We present the task differently, as testing the simple harmonic oscillator model $T=2\pi\sqrt{\frac{L}{\textrm{g}}}$ (Cornell), as testing Galileo’s claims (Tufts), or simply as studying what affects the period (UWB). We guide it differently, providing instructions for students’ experimentation and statistical analyses (UWB and Cornell) or leaving almost everything up to the students (Tufts). The project was funded to study students’ thinking, but as it ended we became interested to study our own: How did we come to design our labs so differently, given the similarities in our perspectives?
@@ I.1 Theoretical frameworks
We were inspired to compare our designs by Boudreaux & Elby [boudreaux_how_2020], who described how their theoretical perspectives shaped their designs of tutorials on static friction. They demonstrated multiple ways in which their distinct theoretical dispositions gave rise to differences in tutorials covering the same content. Boudreaux worked from the “difficulties framework,” which attributes “a degree of stability in student conceptions”: “The strategy of challenging and de-settling non-normative ideas relies on those ideas maintaining coherence long enough to be held up, examined, and modified by the learner as they progress through the curriculum.” This perspective supported designing a tutorial to “elicit, confront, and resolve” incorrect conceptions.
@@
Elby worked from “the resources framework,” according to which “students’ responses to physics questions do not always correspond to stable conceptions.” Rather, patterns in students’ thinking depend on the context, “with contextual cues tipping students” to different ways of thinking. This perspective supported designing a tutorial to manipulate contextual cues that could tip students into activating different patterns of resources, and guiding them to notice and work to reconcile the differences.
@@
The authors were clear that explicit employment of theoretical frameworks did not fully determine the tutorial designs. They also drew on their intuitions and experiences, which they note are deeply entangled with their own theoretical dispositions: “theoretical orientations and instructional intuitions are not disjoint streams of knowledge that independently inform curriculum writing, but rather that they interact as they develop over time.”
In contrast to Boudreaux & Elby, we see ourselves as starting from the same theoretical perspectives. Like Elby, we all take a resource-based view of student knowledge, reasoning, and framing [hammer_resources_2005]. We see student thinking as sensitive to context in complex ways. In our prior work we have considered dynamics at multiple scales of time and system, from individual minds in specific moments, to small collaborative groups over semesters, to classroom and school communities (e.g., [scherr_student_2009, Frank_interactional_2012, radoff_its_2019, dini_case_2017, holmes_toolboxes_2017, quinn_group_2020]).
Rather than presume stability, we expect the dynamics of student thinking can be significantly influenced by how tasks are presented (e.g., [Goodhew_resources_2021, Ford_game_2005, Frank_interactional_2012]), instructor or TA interactions (e.g., [sundstrom_instructing_2023, wu_instructor_2022, Scherr_tainteractions]), intragroup social dynamics (e.g., [conlin_making_2018, zhang_navigating_2021, scherr_student_2009, watkins_positioning_2018]), and practices established in the classroom community (e.g., [appleby_disciplinary_2021, gouvea_motivating_2022]). A resource-based perspective does not rule out stability, however. The theoretical idea of “resources” is of stable aspects of knowledge that can be active or not, depending on the context. Some resources bear on physical mechanism, such as a basic sense that a stronger cause will produce a stronger effect, or that motion “dies away” if there is nothing to sustain it [disessa_1993].
@@
Some resources bear on kinds of activities, such as “storytelling” or “theft.” These are often referred to as “frames” [tannen1993framing]. Our work has documented how student can frame lab as a matter of “jumping through hoops” [phillips_not_2021], and/or expect that success means confirming a known result [smith_how_2020, descamps_dynamics_2024]. These are laboratory versions of “doing school” [jimenez-aleixandre_doing_2000]: following “cookbook” instructions to arrive at the intended results, the findings expert authorities have determined to be correct. Unfortunately, these expectations are often at odds with the practices and values of empirical science.
I.2 Similarities and differences across our labs
@@
In designing our labs, we all anticipate that many students will arrive with expectations they have formed over years of experience in traditional labs, disposed to frame experimental work as a matter of following instructions to arrive at the correct (predetermined) conclusions [smith_how_2020]. Part of what we all hope to accomplish with the pendulum lab is to disrupt that framing and to promote students’ seeing lab as a place for genuine inquiry, such as to make sense of a surprising phenomenon [smith_surprise_2018]. We expect students have resources for the beginnings of empirical science. Given the chance, students can and will be curious about the world, interested and able to try things to learn more about phenomena they notice.
@@
The instructional materials for the Cornell curriculum can be found on PhysPort, on QUBES CourseSource for the UWB curriculum [QUBES], and the Tufts curriculum is in the process of being As would be expected given our similar views and goals, our labs are similar in several ways. We all see the main purpose of introductory labs as helping students to develop abilities in empirical science, rather than to reinforce concepts from lecture [holmes_value_2017, smith_best_2021]. The instructions, assessment, and feedback we provide emphasize practices of empirical science and challenge expectations that success means arriving at a “correct” answer. They give students time and autonomy to design, test, and refine their experimental methods and analyses, and all offer them significant latitude to draw and defend their own conclusions. Our designs all depend to some extent on “responsive teaching” [ResponsiveTeaching] by the instructors, that is, close attention to and engagement with student thinking. Instructors’ engagement with student thinking is not limited to individual or small group interactions, as each of our curricula builds in several opportunities for classroom community-level idea sharing and discussions.
@@
Still, as we noted above and will elaborate below, our labs differ in several, significant ways. We decided to look into those differences, and in what follows we report on what we have found, that variations in our designs arise from (1) differences in our expectations of students at our respective institutions, (2) variations in our ancillary pedagogical goals, and (3) nuances of our theoretical perspectives. We hope that our comparisons provide insight into and concrete examples of how learning theories guide curriculum development. We will try for transparency in our reasoning, to expose uncertainties and flexibilities, to argue for expectations of complexity, and to motivate views of instruction and instructional design as ongoing and dynamic.
I.3 The plan for this article
@@
In the next section we review the literature on the design of “non-traditional” instructional labs, with a specific emphasis on their theoretical bases, both implicit and explicit. Then we move on to our curricula: first, in sec. III, we discuss the institutional and instructional context of the curriculum development; then, in sec. IV, we describe our respective approaches and their ongoing evolution in turn and in some detail; in sec. V we compare these approaches. In sec. VI we step back to make sense and discuss the implications of our comparison before, finally, concluding in sec. VII.
@@ II Literature Review: Lab Curriculum Research and Development
Physicists, educators, and education researchers have long expressed dissatisfaction with traditional laboratory instruction [otero_past_2017, smith_best_2021, reif_teaching_1979, thornton_tools_1987, may_historical_2023]. We understand “traditional” labs to be defined by i) aims to reinforce content knowledge ii) activities that demonstrate or verify physical concepts and iii) prescriptive experimental protocols. Despite their goal of supporting conceptual learning, evidence indicates they are generally unsuccessful at it [thornton_tools_1987, etkina_design_2010, holmes_toolboxes_2017, smith_direct_2020]. Research also indicates that traditional labs negatively impact the development of expertlike epistemological attitudes and beliefs. [wilcox_impact_2016, wilcox_open-ended_2016, wilcox_students_2017, walsh_skills-focused_2022].
Reshaping laboratory instruction with an eye towards authentic disciplinary and experimental practices has notable historical precedent [otero_past_2017, may_historical_2023]. A number of research projects have developed and documented alternative curricula, often inspired by educational theory:
  • RealTime Physics (RTP; [sokoloff_realtime_2007, thornton_tools_1987, thornton_learning_1990])
  • the Investigative Science Learning Environment (ISLE; [etkina_design_2010, brookes_implementing_2020]);
  • Thinking Critically in Physics labs, first at the University of British Columbia [holmes_doing_2013, holmes_making_2014, holmes_impact_2015] and continuing at Cornell [holmes_operationalizing_2019, kalender_restructuring_2021, phillips_not_2021, sundstrom_instructing_2023];
  • the instructional labs at CU-Boulder (e.g., [lewandowski_initial_2018, werth_impacts_2022]
  • the Design, Analysis, Tools, and Apprenticeship lab (DATA lab; [funkhouser_design_2019]).
Some of these learning environments have documented evidence that they support productive experimental activity both in quantity and quality [etkina_design_2010, holmes_developing_2020] and improve students’ epistemological attitudes and beliefs [wilcox_open-ended_2016, smith_direct_2020, brewe_modeling_2009]. At the same time, to generalize the motivations of these curricula as engagement in genuine experimental inquiry or skills-focused designs loses the specificity necessary for researchers to rigorously evaluate and compare them and, crucially, for educators to effectively adapt and implement them. These lab curricula vary in their specific goals, theoretical influences, and pedagogical features and structures. In the next section, we discuss these curricula in more detail by focusing on the differing roles of theory in their development and documentation.
@@ II.1.1 Semi-Empirical
The DATA and CU-Boulder lab curricula are presented as the conclusion of a largely empirical transformation process. In each case, the authors interviewed interested faculty about their goals for laboratory education, synthesized the themes of those interviews into consensus goals, and then iteratively refined the goals along with faculty members. These consensus goals provide the basis for the design of the curriculum.
For the DATA lab, they settled on four learning goals: experimental process, data analysis, collaboration, and communication. Notably, the “design was divorced from the specific physics content.” This feature was explicitly motivated by their empirical findings rather than by theoretical considerations: “the learning goals developed from a faculty consensus design did not include specific content” [funkhouser_design_2019]. To select and design instructional materials and course structures, Funkhouser et al. applied Constructive Alignment [biggs_alignment_1996]: the alignment of an instructional system based on constructivist principles. In this scheme, course objectives should articulate how students can demonstrate understanding through actions, students should be put in situations that can elicit such “performances of understanding,” and students should be assessed based on their performances.
The DATA lab is a stand-alone course and two semester sequence. The typical lab activity takes place over two weeks and students generally follow the same two-week trajectory: the pre-lab homework informs the qualitative exploration and prediction phase and the first week usually culminates with students’ designing their experimental procedure; in the second week, students conduct their investigation and iteratively improve their procedure as they analyze their data and generate interpretable results.
Details of the initial CU-Boulder lab curriculum, also a stand-alone lab course, can be found in a series of PERC papers [lewandowski_student_2018, pollard_impact_2018, pollard_transforming_2018, lewandowski_initial_2018]. The design of this curriculum was motivated by five main learning goals:
@@
  • students should demonstrate “set-like” reasoning about data
Additionally, in other work (e.g., [werth_impacts_2022]), the CU-Boulder team refers to their initial lab as seeking to develop skills. At the same time, the research teams notes that, due to various constraints, the course retained many features of a traditional lab [pollard_impact_2020].
Nevertheless, Pollard et al. [pollard_impact_2020] highlight several transformed features they expect to be salient for student learning: for a pre-lab assignment, students watch short, interactive pre-lab videos; the experimental activities require students to measure a previously unknown quantity or outcome, as opposed to verifying values learned in lecture or listed in a textbook; students are often asked to use these measurements to make a prediction related to their experiment and then, once they measure that prediction, compare results with other groups—which is expected to support the use of and reflection on uncertainty values.
Although presented as a mostly empirical process, we contend that there is non-trivial theoretical work involved in the development and design of these curricula. For example, a notable feature of the DATA lab is that, for a given activity, lab groups each conduct a different experiment and there is no set end point for these experiments. Funkhouser et al. mention that this feature “has added benefits for student’s ownership and agency of the work.” It could be that ownership and agency are understood by the designers to support productive engagement and thus were part of the decision-making process; another possibility is that increased ownership and agency are serendipitous benefits afforded by this choice, which was influenced by other considerations.
The work of specifying how to accomplish learning goals can be understood as a process of making hypotheses or conjectures about how to best support learning [sandoval2014, Cobb2007]. In this way, this empirical process obscures the justification of design decisions. Instead, curricular features are presented as naturally emerging from learning goals. In doing so, the documentations of course transformations do not provide actionable insight for educators selecting among alternatives as they adapt design features to different contexts. Accounts of the reasoning that motivates design decisions in a curriculum are especially important given our understanding of the variation in how educators take up research-based instructional methods [dancy_how_2016, turpen_construction_2010, affriyenni_navigating] and the consequential role of the broader (activity) system within which specific design features are implemented [tasar_characterizing_2023].

### v2

@@ Abstract
We present three versions of a pendulum lab activity to explore how theoretical commitments, motivations, and aspirations reflect in curriculum design. In earlier work, Boudreaux & Elby [7] discussed how different theoretical perspectives led them to different designs of tutorials. Here, we discuss our finding that even starting from the same theoretical perspective and having similar goals for students can lead to differences of design. We give three interacting reasons why our labs diverge: We have different expectations of students at our respective institutions; we have different ancillary goals; and we draw different implications from our shared theoretical commitments. Our account demonstrates the complexity of the relationship between theory, goals, and curriculum design. In this way, it adds to prior arguments for the importance of designers’ articulating the reasoning for their choices as well as for the possible value of instructors’ responsive adaptations to curricula.
@@ I Introduction
For longer than any of the authors of this paper have been learning physics, introductory labs have included a pendulum activity. Often, a goal of the activity is for students to demonstrate that the period of a pendulum has a small but measurable dependence on the amplitude of its swing, a result at odds with the simple harmonic oscillator model and Galileo’s claim in Two New Sciences [23]. Such a lab provides a context for learning practices and values of empirical science: to see the dependence, students need both to take care in their measurements and to be aware of the precision they have obtained. [^1] They do not need sophisticated technology! Mechanical stopwatches work fine.
Each of us has assigned pendulum labs at the start of our introductory courses, at our respective institutions, with goals of introducing practices and values of empirical science. We want the activities to promote students’ attention to experimental error as well as reflection and conversation about how science produces and assesses knowledge. These goals reflect our shared view of students: that they have mostly experienced step-by-step instructions for experiments that demonstrate or validate theoretical claims from lecture, and need support to engage in more deeply scientific practices.
The similarities in our interests, goals, and views motivated research on how students may shift their approach to labs, to be less about “doing school” and more about “doing physics” [77, 79, 76]. That is, we have asked: when and how do students come to frame introductory labs in ways that support progress toward practices and values of empirical science? We were supported by the National Science Foundation for three years to collect and study data from our courses. [^2] NSF DUE Awards: 2000394, 2000739, 2000711. We paid particular attention to the pendulum labs, because they were the students’ first engagement with activities different from their expectations. The project led to several publications analyzing student framings, including with respect to confirmation bias and epistemic agency [82, 57, 44, 83, 14].
Throughout, we authors were aware of differences in our lab designs. We present the task differently: as testing the simple harmonic oscillator model $T=2\pi\sqrt{\frac{L}{\textrm{g}}}$ (Cornell), as testing Galileo’s claims (Tufts), or simply as studying what affects the period (UWB). We guide it differently, providing instructions for students’ experimentation and statistical analyses (UWB and Cornell) or leaving almost everything up to the students (Tufts). The project was funded to study students’ thinking, but as it ended we became interested to study our own: How did we come to design our labs so differently, given the similarities in our perspectives?
@@ I.1 Theoretical frameworks
We were inspired by Boudreaux & Elby [7], who described how their theoretical perspectives shaped their designs of tutorials on static friction. They demonstrated multiple ways in which their distinct theoretical orientations gave rise to differences in tutorials covering the same content. Boudreaux worked from a “difficulties framework,” which attributes “a degree of stability in student conceptions”: “The strategy of challenging and de-settling non-normative ideas relies on those ideas maintaining coherence long enough to be held up, examined, and modified by the learner as they progress through the curriculum.” This perspective supported designing a tutorial to “elicit, confront, and resolve” incorrect conceptions.
@@
Elby worked from a "resources framework,” according to which “students’ responses to physics questions do not always correspond to stable conceptions.” Rather, patterns in students’ thinking depend on the context, “with contextual cues tipping students” to different ways of thinking. This perspective supported designing a tutorial to manipulate contextual cues that could tip students into activating different patterns of resources and guide them to notice and work to reconcile the differences.
@@
The authors were clear that explicit employment of theoretical frameworks did not fully determine their tutorial designs. They also drew on their intuitions and experiences, which they note are deeply entangled with their own theoretical dispositions: “theoretical orientations and instructional intuitions are not disjoint streams of knowledge that independently inform curriculum writing, but rather that they interact as they develop over time.”
I.2 Our theoretical orientations
In contrast to Boudreaux & Elby, we see ourselves as starting from the same theoretical perspective. Like Elby, we all take a resource-based view of student knowledge, reasoning, and framing [28]. The theoretical idea of “resources” is of “pieces” [17] of knowledge that can be active or not, depending on the context. Some resources bear on physical mechanism, such as a basic sense that a stronger cause will produce a stronger effect, or that motion “dies away” if there is nothing to sustain it [17]. Some bear on kinds of activities and situations, influencing which features people recognize and pay attention to, as well as their judgments about appropriate actions. These resources are often referred to as “frames” and their activation as "framing" [84, 28, 71].
A resources perspective does not presume stability. Resources may become active or not depending on the context. For example, in the context of cold weather, it is easy and intuitive to think “gloves keep hands warm,” while in the context of baking, “gloves keep hands cool.” Much of the work of physics is in connecting reasoning in contexts that start out as separate. Similarly, for example, someone framing what is taking place as following instructions may not notice or think to check whether the instructions make sense to them, as they would if they framed the moment instead as their participating in a decision.
Taking this perspective, we expect the dynamics of student thinking to be significantly influenced by how tasks are presented (e.g., [25, 20, 21]), instructor or TA interactions (e.g., [83, 95, 73]), intragroup social dynamics (e.g., [11, 96, 71, 89]), and practices established in the classroom community (e.g., [3, 26]). In our prior work, we have considered dynamics at multiple scales of time and system, from individual minds in specific moments, to small collaborative groups over semesters, to classroom and school communities (e.g., [71, 21, 62, 15, 38, 61]).
@@ I.2 Our theoretical orientations
Our work in instructional labs has documented how students can frame lab as a matter of “jumping through hoops” [57], and/or expect that success means confirming a known result [79, 14]. These are laboratory versions of “doing school” [46]: following “cookbook” instructions to arrive at the intended results, the findings expert authorities have determined to be correct. Unfortunately, these expectations are often at odds with the practices and values of empirical science.
The frames highlighted above, and students’ framing in general, have profound consequences on the nature and quality of students’ participation in labs. That is to say, how students frame the lab shapes the ways and extent to which they take up epistemic agency. [83, 14, 57, 47] Epistemic agency is the prerogative for individuals or groups to participate in “the whole range of components of knowledge building goals, strategies, resources, evaluation of results, and so on” [70] (p. 108). In our research and in our designs, epistemic agency is central to our understanding of what constitutes productive participation in scientific knowledge construction.
I.3 Similarities and differences across our labs
@@
In designing our labs, we all anticipate that many students will arrive with expectations they have formed over years of experience in traditional labs, disposed to frame experimental work as a matter of following instructions to arrive at the correct (predetermined) conclusions [79]. Part of what we all hope to accomplish with the pendulum lab is to disrupt that framing and to promote students’ seeing lab as a place for genuine inquiry, such as to make sense of a surprising phenomenon [77]. We expect students have resources for the beginnings of empirical science. Given the chance, students can and will be curious about the world, interested and able to try things to learn more about phenomena they notice.
@@
The instructional materials for the Cornell curriculum can be found on PhysPort; instructional materials for the UWB curriculum can be found on QUBES CourseSource [72]; instructional materials for the Tufts curriculum are attached as supplemental materials. As would be expected given our similar views and goals, our labs are similar in several ways. We all see the main purpose of introductory labs as helping students develop abilities in empirical science, rather than reinforcing concepts from lecture [39, 78]. The instructions, assessment, and feedback we provide emphasize practices of empirical science and challenge expectations that success means arriving at a “correct” answer. They give students time and autonomy to design, test, and refine their experimental methods and analyses, and all offer them significant latitude to draw and defend their own conclusions. Finally, all of our designs depend to some extent on “responsive teaching” [67], that is the lab instructors’ close attention to and engagement with student thinking, in individual and small group interactions, as well as in classroom community-level idea sharing and discussions.
@@
Still, as we noted above and will elaborate below, our labs differ in several significant ways. We decided to look into those differences, and in what follows we report on what we have found: Variations in our designs arise from (1) differences in our expectations of students at our respective institutions, (2) variations in our ancillary pedagogical goals, and (3) nuances of our theoretical perspectives. We will try to make our reasoning transparent, to expose uncertainties and flexibilities in our designs, to argue for expectations of complexity, and to motivate views of instruction and instructional design as ongoing and dynamic.
We hope that our comparisons provide insight into and concrete examples of how learning theories guide curriculum development. Scholars have long argued, and empirical evidence indicates, that it is important for teachers implementing curriculum to understand the purposes and aims of its design [74, 75, 5, 13, 30, 18, 65]. Our comparative analysis supports these arguments, exposing "hidden decisions" [30] and "pedagogical judgments" [13] that led us to different versions of the pendulum lab.
@@
In the next section, we review the literature on the design of “non-traditional” instructional labs, with an emphasis on their theoretical bases, both implicit and explicit. Then we move on to our curricula: first, in sec. III, we discuss each of the institutional and instructional contexts; in sec. IV, we describe our respective approaches and their ongoing evolution in turn and in some detail; and in sec. V we compare these approaches. In sec. VI we step back to make sense and discuss the implications of our comparison before, finally, concluding in sec. VII.
@@ II Literature Review: Lab Curriculum Research and Development
Physicists, educators, and education researchers have long expressed dissatisfaction with traditional laboratory instruction [55, 78, 63, 86, 52]. We understand “traditional” labs to be defined by i) aims to reinforce content knowledge, ii) activities that demonstrate or verify physical concepts, and iii) prescriptive experimental protocols. Despite such labs’ goal of supporting conceptual learning, evidence indicates they are generally unsuccessful at it [86, 19, 38, 80]. Research also indicates that traditional labs negatively impact the development of expertlike epistemological attitudes and beliefs [92, 94, 93, 88].
Reshaping laboratory instruction with an eye towards authentic disciplinary and experimental practices has notable historical precedent [55, 52]. A number of research projects have developed and documented alternative curricula, often inspired by educational theory:
  • RealTime Physics (RTP; [81, 86, 85])
  • Investigative Science Learning Environment (ISLE; [19, 9]);
  • Thinking Critically in Physics labs, first at the University of British Columbia [32, 33, 34] and continuing at Cornell [37, 47, 57, 83];
  • the instructional labs at CU-Boulder (e.g., [50, 91])
  • the Design, Analysis, Tools, and Apprenticeship lab (DATA lab; [22]).
Some of these learning environments have documented evidence that they support productive experimental activity in both quantity and quality [19, 35] and improve students’ epistemological attitudes and beliefs [94, 80, 8]. At the same time, generalizing the motivations of these curricula as engagement in genuine experimental inquiry or skills-focused designs loses the specificity necessary for researchers to rigorously evaluate and compare them and, crucially, for educators to effectively adapt and implement them. These lab curricula vary in their specific goals, theoretical influences, and pedagogical features and structures. In the next section, we discuss these curricula in more detail by focusing on the differing roles of theory in their development and documentation.
@@ II.1.1 Semi-Empirical
The DATA and CU-Boulder lab curricula are presented as the conclusion of a largely empirical transformation process. In each case, the lab developers interviewed interested faculty about their goals for laboratory education, synthesized the themes of those interviews into consensus goals, and then iteratively refined the goals along with faculty members. These consensus goals provided the basis for the design of the curriculum. We refer to this process as semi-empirical because, while the overall course transformations incorporated empirical research, the design decisions were not explicitly grounded in empirical (or theoretical) explanations.
For the DATA lab, they settled on four learning goals: experimental process, data analysis, collaboration, and communication. Notably, the “design was divorced from the specific physics content.” This feature was explicitly motivated by their empirical findings rather than by theoretical considerations: “the learning goals developed from a faculty consensus design did not include specific content” [22]. To select and design instructional materials and course structures, Funkhouser et al. applied Constructive Alignment [6]: the alignment of an instructional system based on constructivist principles. In this scheme, course objectives should articulate how students can demonstrate understanding through actions; students should be put in situations that can elicit such “performances of understanding,” and students should be assessed based on their performance.
The DATA lab is a stand-alone course and two semester sequence. The typical lab activity takes place over two weeks and students generally follow the same two-week trajectory: pre-lab homework informs the qualitative exploration and prediction phase, and the first week culminates in students’ designing their experimental procedure; in the second week, students conduct their investigation and iteratively improve their procedure as they analyze their data and generate interpretable results.
Details of the initial CU-Boulder lab curriculum, also a stand-alone lab course, can be found in a series of PERC papers [49, 58, 59, 50]. The design of this curriculum was motivated by five main learning goals:
@@
  • students should demonstrate “set-like” reasoning about data.
Additionally, in other work [91], the CU-Boulder team refers to their initial lab as seeking to develop skills. The research team also notes that, due to various constraints, the course retained many features of a traditional lab [60].
Nevertheless, Pollard et al. [60] highlight several transformed features they expect to be salient for student learning: for a pre-lab assignment, students watch short, interactive videos; the experimental activities require students to measure a previously unknown quantity or outcome, as opposed to verifying values learned in lecture or listed in a textbook; students use these measurements to make a prediction related to their experiment and then, once they measure that prediction, compare results with other groups. That comparison is expected to support the intentional use of uncertainty values. Additionally, most of the lecture sessions (which are infrequent) are dedicated to measurement uncertainty concepts.
Although presented as a mostly empirical process, we contend that there is non-trivial theoretical work involved in the development and design of these curricula. For example, a feature of the DATA lab is that, for a given activity, each lab group conducts a different experiment, with no set end point. Funkhouser et al. mention that this feature “has added benefits for [each] student’s ownership and agency of the work.” It could be that ownership and agency were understood by the designers to support productive engagement and thus were part of the decision-making process. Another possibility is that increased ownership and agency were serendipitous benefits afforded by this choice, which was influenced by other considerations.
This approach obscures the justification of design decisions. Curricular features are instead presented as naturally emerging from learning goals. As a result, the documentation of curriculum transformations does not provide actionable insight for educators enacting the curriculum or selecting among alternatives as they adapt design features. The work of specifying how to accomplish learning goals can be productively understood as a theory-laden process of making hypotheses or conjectures about how to best support learning [69, 10]. Accounts of the theory-laden reasoning that motivates design decisions in a curriculum are especially important given our understanding of the variation in how educators take up research-based instructional methods [12, 87, 1], the consequential role of the broader (activity) system within which specific design features are implemented [97, 29], and the ways in which an awareness of this reasoning can support responsive instructional moves [65]